OpenStudy (antaniece):

What is the solutions to the follow systems ? Y-3=x X^2-6x+13=y A. (-5,2) B. (-2,1) C. (2,5) D. (8,5)

OpenStudy (latinc):

first, you know what x is so go ahead and input your x value below

OpenStudy (scooby6363):

(2,5) Solve for y in both equations. The second equation cannot be factored y = x + 3 y = x^2 - 6x + 13 At this moment you can graph both functions and view at which points do the two functions meet. I got two solutions, (2,5) and (5,8). Since there is no (5,8) you should go with (2,5)

OpenStudy (latinc):

----> x^2 -6x+13=x+3